OpenStudy (anonymous):

What should be measured to calculate the average rate at which water flows in a dam? Distance of water flow and duration of water flow Number of liters of water flowing and duration of water flow Distance of water flow and size of the dam from which water flows Number of liters of water flowing and size of the dam from which water flows

random231 (random231):

number of liters of water and duration of flow

OpenStudy (anonymous):

I'd say the second answer because you need an amount and a time frame

OpenStudy (anonymous):

that was my first guess, thanks

random231 (random231):

cool

OpenStudy (anonymous):

wanna help me with another question ?

random231 (random231):

sure

OpenStudy (anonymous):

The equation below shows the perimeter f(x) of several rectangular pieces of cardboard having width x inches: f(x) = 52 + 2x The perimeter of each piece is less than 100 inches. Which of the following best describes the domain of f(x)? All real numbers greater than 0 and less than 100 All real numbers greater than 0 and less than 24 All integers greater than 0 and less than 100 All integers greater than 0 and less than 24

random231 (random231):

the first one

OpenStudy (anonymous):

If you set up an equation like this\[100>52+2x\] and then just solve for the equation.

OpenStudy (anonymous):

No it's not the first one.

random231 (random231):

why not @TheBadWolf ?

OpenStudy (anonymous):

It's the second, because if you put a number into that equation that is greater than 24 your function is over 100

OpenStudy (anonymous):

so its All real numbers greater than 0 and less than 24

random231 (random231):

no wait

OpenStudy (anonymous):

yeah

random231 (random231):

in the question its given each piece is less than 100 therefore x can attain values between 0-100

random231 (random231):

u r wrong @TheBadWolf

OpenStudy (anonymous):

omg guys !

random231 (random231):

yeah f(x) IS THE VALUE OF THE PERIMETER OF THE JOINED PIECES AND NO CONDITION IS GIVEN THERE THAT IT CANT BE GREATER THAN 100. ITS THE VALUE OF EACH PIECE THAT CANNOT BE GREATER THAN 100.

OpenStudy (anonymous):

THe function \[f \left( x \right)= 52-2x\] has to be less then 100

random231 (random231):

NO

OpenStudy (anonymous):

ok im on another level of confusion now

random231 (random231):

just a second @yeah_i_suck_at_math . i hav a doubt with the question

random231 (random231):

is it perimeter of each piece is given by f(x) or all the pieces?

OpenStudy (anonymous):

I think the 52 = 2L and the 2x symbolizes 2w. when you add them together you get the perimeter

OpenStudy (anonymous):

so in order for the perimeter to be less than 100 the function has to be less than 100

random231 (random231):

@TheBadWolf what u say is right. But im having a confusion. as the question says "the perimeter of several pieces of cardboard is given by f(x)" it says several. so is f(x) the perimeter of several pieces or of 1 piece only?

OpenStudy (anonymous):

so ....

OpenStudy (anonymous):

all of the perimeters are less than 100 and all the lengths are the same. the onlything that changes is the width. the range for the width while keeping the perimeter under 100 is All real numbers greater than 0 and less than 24

OpenStudy (anonymous):

so it was B all along ?

random231 (random231):

okay i admit i was wrong. sorry @yeah_i_suck_at_math nd @TheBadWolf

random231 (random231):

yeah

OpenStudy (anonymous):

its ok

OpenStudy (anonymous):

help me with another one ? it looks simple..

random231 (random231):

sure

OpenStudy (anonymous):

The weights of four packets of sugar are shown below: Packet A: 4.35 pounds Packet B: 4.42 pounds Packet C: 4.48 pounds Packet D: 4.39 pounds Sandy has a scale that measures to the nearest tenth of a pound. If Sandy weighs each packet on her scale, the heaviest reading will be _____ pounds.

OpenStudy (anonymous):

Btw no hard feelings @random231 it's all good. This question wants you to rund to the nearest tenth. The tenth spot is the one directly to the right of the decimal. Packet A would read as 4.4 Packet B would also read as 4.4 Packet C would read as 4.5 and Packet D would again read as 4.4 So packet D is the heaviest reading the scale will register

OpenStudy (anonymous):

packet d is 4.4 and so is a and b ...

OpenStudy (anonymous):

i think u meant c is th heaviest

OpenStudy (anonymous):

yeah sorry whoops

OpenStudy (anonymous):

do i enter it as 4.48

OpenStudy (anonymous):

or 4.5

OpenStudy (anonymous):

doesnt say to round

random231 (random231):

write 4.48

OpenStudy (anonymous):

Sandy has a scale that measures to the nearest tenth of a pound...

OpenStudy (anonymous):

well the scale automatically estimates because it only shows the first decimal place. The highest READING it will give is 4.5 and that would fill in your blank

OpenStudy (anonymous):

help me with another one ?

OpenStudy (anonymous):

Jack wants to fill a rectangular box with sand. The length of the sand box is 3 feet, width is 6 inches, and height is 2.4 inches. Each bag of sand contains 0.15 cubic foot of sand. How many bags of sand will Jack need to fill the box completely? [1 foot = 12 inches]

random231 (random231):

volume of sand required = 3*0.5*0.2 cubic feet

OpenStudy (anonymous):

well to find the volume you multiply the hieght, width, and length so\[36 \times 6 \times 2.4=\] I use 36 because 3 ft is equal to 36 inches

OpenStudy (anonymous):

when you find the answer to that equation you divide it by 12 to turn it back to feet and then by .15 to figure out how many bags you need

random231 (random231):

so no of bags required = (3*0.5*0.2)/0.15

OpenStudy (anonymous):

so 518.4

OpenStudy (anonymous):

I got 288 bags

OpenStudy (anonymous):

288 bags ?

OpenStudy (anonymous):

yea so its 288

random231 (random231):

288?

OpenStudy (anonymous):

yeah... ?

OpenStudy (anonymous):

No actually I'm wrong, because 36x6x2.4=518.4in cubed but we need feet cubed so divide by 12 three times because it needs to be cubed

random231 (random231):

ans = 2bags

OpenStudy (anonymous):

and then dived by.15 and random is correct

random231 (random231):

why dont u convert inches to feet in beginning?

random231 (random231):

6=0.5 ft 2.4=0.2ft

OpenStudy (anonymous):

so 2 bags is the answer ?

random231 (random231):

yea

OpenStudy (anonymous):

it was easier for me I guess. In hindsight that works better, but yes 2 bags is your answer

OpenStudy (anonymous):

ok how about this one ? thanks btw guys A car can travel 25 miles per gallon. Approximately how many gallons of fuel will the car need to travel 30 km? [1 mile = 1.6 km] 0.05 0.52 0.75 0.83

random231 (random231):